### v1

@@ 6.1 Effectiveness of Low-Level Feature Extractors
Since each domain has proposed specific feature extractors, to explore their effectiveness under a unified domain, we conduct experiments using 6 backbones combined with 4 different extractors in the shallow layer under the aforementioned IFF-Protocol setting. The extractors are BayarConv [3] for noise aritfacts, Sobel [5] for edge arifacts, DCT [1] and FPH [43] for frequency artifacts. Details of each extractor can be found in Appendix G.1.
Results in Table 8 show the AUC differences between versions of each backbone using the four different feature extractors and those without, averaged across all test datasets for each task. All backbones except EfficientNet [55] show performance drops after using feature extractors, indicating that under the IFF-Protocol, where training data includes sufficient manipulation types and image quantity, models do not rely on the additional information provided by feature extractors. However, due to its lightweight nature, EfficientNet still benefits from the use of feature extractors. The results suggest that feature extractors may only be beneficial for detection on small-scale datasets, with limited manipulation types, or when using lightweight models. Details of each domain test datasets AUC scores can be found in Appendix G.2.
@@
Task Extractor = Sobel [5] Extractor = Bayar[3]
@@
Extractor = FPH [43] Extractor = DCT [1]
@@ 6.2 Transferability of Task-Specific Detectors
6.2.1 Cross-Evaluation Between IMDL and Document Benchmarks
Current Document-level detectors’ input–output formats are fully compatible with those of Image Manipulation Detection and Localization models. Leveraging this consistency, we perform a bidirectional evaluation: IMDL detectors are tested on the Document benchmark, and Document detectors are tested on the IMDL benchmark. This cross-testing enlarges the effective model pool for each benchmark and allows us to probe detector generality beyond their original task scopes.
@@ 6.2.1 Cross-Evaluation Between IMDL and Document Benchmarks
Table 4 reports the within-domain results obtained under the original Document benchmark split, whereas Table 5 presents the cross-domain scores produced by our newly introduced generalization protocol. Across both settings, IMDL detectors demonstrate strong competitiveness in the document forensics scenario. In the conventional split, the Cat-Net [22, 43, 6] family achieves the best average F1, confirming the merit of its hierarchical “cat-net” paradigm. Under the more challenging cross-domain evaluation, PSCC-Net [29] displays markedly better generalization, suggesting that progressive spatial modeling captures cues for document manipulation localization. We expect future work to further investigate the underlying mechanisms behind PSCC-Net.
@@
Following the MVSS training protocol [36], all Document-oriented models are trained on the CASIAv2 dataset [11] and evaluated on five standard IMDL test sets. As shown in Table 9, the dual-branch architecture of CAFTB [52] achieves the best overall performance among Document models when transferred to IMDL tasks—an outcome that aligns with the design philosophy of the current SoTA model Mesorch [70], which also emphasizes dual-branch learning.
@@
CAFTB [52](TOMM24) 0.6234 0.2557 0.6731 0.3526 0.3770 0.4564
DTD [43](CVPR23) 0.3535 0.1482 0.6470 0.1749 0.1811 0.3009
FFDN [6](ECCV24) 0.5012 0.2670 0.6085 0.2516 0.2957 0.3848
TIFDM [12](TCE24) 0.3675 0.2087 0.3572 0.1634 0.2333 0.2660
@@ 6.2.2 Extending IMDL Detectors to AIGC and Deepfake Benchmarks
We fine-tune representative IMDL detectors on the training split of the AIGC benchmark and report cross-generator performance in Table 3. The training settings and other configurations are consistent with those used in the previously mentioned AIGC benchmark setup. The results show that techniques from IMDL, such as noise print (TruFor) and multiscale analysis (IML-ViT), remain effective for AIGC detection.
@@
We train each IMDL detector on the FF++-c23 training split and evaluate on all remaining deep-fake test sets; the scores are given in Table 10. When compared to all baselines in the deepfakebench [67], Cat-Net attains the best performance in the within-domain setting, while Mesorch achieves the highest average accuracy in the cross-domain evaluation, establishing new state-of-the-art results in both regimes.
@@
FF-c23 FF-c40 FF-DF FF-F2F FF-FS FF-NT Avg. CDFv1 CDFv2 DFD DFDC DFDCP Fsh UADFV Avg.
CAT-Net [22](IJCV22) 0.9855 0.8351 0.9946 0.9891 0.9905 0.9689 0.9606 0.7489 0.7511 0.8086 0.7196 0.7043 0.6626 0.9282 0.7605
IML-ViT [35](Arxiv) 0.9614 0.8483 0.9810 0.9728 0.9793 0.9206 0.9439 0.7250 0.7419 0.7770 0.7516 0.7237 0.6149 0.8718 0.7437
Mesorch [70](AAAI25) 0.9815 0.8445 0.9889 0.9890 0.9859 0.9636 0.9589 0.7900 0.7649 0.8468 0.7806 0.7345 0.6638 0.9623 0.7918
MVSS-Net [5](ICCV21) 0.9723 0.8140 0.9779 0.9865 0.9868 0.9458 0.9472 0.6969 0.6839 0.7749 0.6764 0.6785 0.5841 0.8871 0.7117
PSCC-Net [29](TCSVT22) 0.7147 0.6536 0.7751 0.6414 0.6460 0.6065 0.6729 0.4602 0.5898 0.5565 0.5945 0.5845 0.6574 0.8216 0.6092
Trufor [17](CVPR23) 0.7303 0.6848 0.8087 0.7687 0.7084 0.6326 0.7223 0.6551 0.6115 0.5908 0.6017 0.5749 0.5977 0.9526 0.6549
@@ D.3 Metrics
[…]
@@ E.1 Training Details for AIGC Benchmark Implementation
We resize images to 224×224 and apply only basic data augmentations, including flipping, brightness and contrast adjustment, compression, and Gaussian blur. All models are trained for 20 epochs using a cosine decay learning rate schedule, decreasing from $41e-$ 4 to $51e-$ 5. We use the ImageNet split of DiffusionForensics as the training set. For Synthbuster, we use a fully connected layer as the classifier.
@@ E.2 Training Details for Document Image Manipulation Localization Benchmark Models
Model FF++c40 FF-DF FF-F2F FF-NT FF-FS CDFv1 CDFv2 DFD DFDC DFDCP FaceShifter UADFV Average
@@ Appendix G Details of Experiments
Extractor Model FF++c40 FF-DF FF-F2F FF-NT FF-FS CDFv1 CDFv2 DFD DFDC DFDCP FaceShifter UADFV Mean
@@ G.2 Details for Extractor & Backbone in different tasks
We provide the performance differences of 6 backbones with and without 4 different feature extractors across the 4 domains. The table presents detailed results for each individual test dataset. They are Table 15 for IMDL, Table 16 for AIGC, Tabel 17 for Deepfake, and Table 18 for Document.
@@ Appendix H Computational Resources
The experiments were conducted on three different servers. The first server is equipped with two AMD EPYC 7542 CPUs, 256GB RAM, and 6$\times$×NVIDIA A40 GPUs, which was used for all IFF-related experiments. The remaining experiments were performed on two servers: one with a single AMD EPYC 7542 CPU, 256GB RAM, and 4$\times$×NVIDIA RTX 3090 GPUs, and another with two AMD EPYC 7542 CPUs, 256GB RAM, and 8$\times$×NVIDIA RTX 3090 GPUs.

### v2

@@ 6.1 Effectiveness of Low-Level Feature Extractors
Since each domain has proposed specific feature extractors, to explore their effectiveness under a unified domain, we conduct experiments using 6 backbones combined with 4 different extractors in the shallow layer under the aforementioned IFF-Protocol setting. The extractors are BayarConv Bayar_2018 for noise aritfacts, Sobel MVSS_2021 for edge arifacts, DCT ahmed2006discrete and FPH qu2023doctamper for frequency artifacts. Details of each extractor can be found in Appendix G.1.
Results in Table 8 show the AUC differences between versions of each backbone using the four different feature extractors and those without, averaged across all test datasets for each task. All backbones except EfficientNet tan2019efficientnet show performance drops after using feature extractors, indicating that under the IFF-Protocol, where training data includes sufficient manipulation types and image quantity, models do not rely on the additional information provided by feature extractors. However, due to its lightweight nature, EfficientNet still benefits from the use of feature extractors. The results suggest that feature extractors may only be beneficial for detection on small-scale datasets, with limited manipulation types, or when using lightweight models. Details of each domain test datasets AUC scores can be found in Appendix G.2.
@@
Task Extractor = Sobel MVSS_2021 Extractor = BayarBayar_2018
@@
Extractor = FPH qu2023doctamper Extractor = DCT ahmed2006discrete
@@ 6.2 Transferability of Task-Specific Detectors
6.2.1 Cross‑Evaluation Between IMDL and Document Benchmarks
Current Document-level detectors’ input–output formats are fully compatible with those of Image Manipulation Detection and Localization models. Leveraging this consistency, we perform a bidirectional evaluation: IMDL detectors are tested on the Document benchmark, and Document detectors are tested on the IMDL benchmark. This cross‑testing enlarges the effective model pool for each benchmark and allows us to probe detector generality beyond their original task scopes.
@@ 6.2.1 Cross‑Evaluation Between IMDL and Document Benchmarks
Table 4 reports the within‑domain results obtained under the original Document benchmark split, whereas Table 5 presents the cross‑domain scores produced by our newly introduced generalization protocol. Across both settings, IMDL detectors demonstrate strong competitiveness in the document forensics scenario. In the conventional split, the Cat-Net CAT-Net2022 ; qu2023doctamper ; chen2024ffdn family achieves the best average F1, confirming the merit of its hierarchical “cat‑net” paradigm. Under the more challenging cross‑domain evaluation, PSCC-Net liu2022pscc displays markedly better generalization, suggesting that progressive spatial modeling captures cues for document manipulation localization. We expect future work to further investigate the underlying mechanisms behind PSCC-Net.
@@
Following the MVSS training protocol ma2024imdl , all Document-oriented models are trained on the CASIAv2 dataset CASIA_2013 and evaluated on five standard IMDL test sets. As shown in Table 9, the dual-branch architecture of CAFTB song2025caftb achieves the best overall performance among Document models when transferred to IMDL tasks—an outcome that aligns with the design philosophy of the current SoTA model Mesorch zhu2025mesoscopic , which also emphasizes dual-branch learning.
@@
CAFTB song2025caftb (TOMM24) 0.6234 0.2557 0.6731 0.3526 0.3770 0.4564
DTD qu2023doctamper (CVPR23) 0.3535 0.1482 0.6470 0.1749 0.1811 0.3009
FFDN chen2024ffdn (ECCV24) 0.5012 0.2670 0.6085 0.2516 0.2957 0.3848
TIFDM dong2024tifdm (TCE24) 0.3675 0.2087 0.3572 0.1634 0.2333 0.2660
@@ 6.2.2 Extending IMDL Detectors to AIGC and Deepfake Benchmarks
We fine‑tune representative IMDL detectors on the training split of the AIGC benchmark and report cross‑generator performance in Table 3. The training settings and other configurations are consistent with those used in the previously mentioned AIGC benchmark setup. The results show that techniques from IMDL, such as noise print (TruFor) and multiscale analysis (IML-ViT), remain effective for AIGC detection.
@@
We train each IMDL detector on the FF++‑c23 training split and evaluate on all remaining deep‑fake test sets; the scores are given in Table 10. When compared to all baselines in the deepfakebench yan2023deepfakebench , Cat-Net attains the best performance in the within‑domain setting, while Mesorch achieves the highest average accuracy in the cross‑domain evaluation, establishing new state‑of‑the‑art results in both regimes.
@@
$F$-c23 $F$-c40 $F$-DF $F$-F2F $F$-FS $F$-NT Avg. CDFv1 CDFv2 DFD DFDC DFDCP Fsh UADFV Avg.
CAT-Net CAT-Net2022 (IJCV22) 0.9855 0.8351 0.9946 0.9891 0.9905 0.9689 0.9606 0.7489 0.7511 0.8086 0.7196 0.7043 0.6626 0.9282 0.7605
IML-ViT ma2023iml (Arxiv) 0.9614 0.8483 0.9810 0.9728 0.9793 0.9206 0.9439 0.7250 0.7419 0.7770 0.7516 0.7237 0.6149 0.8718 0.7437
Mesorch zhu2025mesoscopic (AAAI25) 0.9815 0.8445 0.9889 0.9890 0.9859 0.9636 0.9589 0.7900 0.7649 0.8468 0.7806 0.7345 0.6638 0.9623 0.7918
MVSS-Net MVSS_2021 (ICCV21) 0.9723 0.8140 0.9779 0.9865 0.9868 0.9458 0.9472 0.6969 0.6839 0.7749 0.6764 0.6785 0.5841 0.8871 0.7117
PSCC-Net liu2022pscc (TCSVT22) 0.7147 0.6536 0.7751 0.6414 0.6460 0.6065 0.6729 0.4602 0.5898 0.5565 0.5945 0.5845 0.6574 0.8216 0.6092
Trufor trufor2023 (CVPR23) 0.7303 0.6848 0.8087 0.7687 0.7084 0.6326 0.7223 0.6551 0.6115 0.5908 0.6017 0.5749 0.5977 0.9526 0.6549
6.3 Grad-CAM Visualization
We use Grad-CAM to visualize the heatmaps of models from the four domains (Capsule-Net (Deepfake), MVSS-Net (IMDL), UnivFD (AIGC), DTD (Doc)) on datasets from each domain, aiming to explore their attention regions, as shown in Figure 2. We use Grad-CAM to visualize the heatmaps of models from the four domains. Models from different domains show both common and distinct attention patterns. For Deepfake, CapsuleNet focuses on specific facial features, while MVSS-Net attends to larger areas. For Doc, CapsuleNet, MVSS-Net, and UnivFD capture overall tampered regions, whereas DTD targets subtle traces like edges and curves of characters.
Figure 2: Grad-CAM visualization (zoomed in for better visualization).
@@ 7 Conclusion
8 Acknowledgment
This research was supported by the Sichuan Province Major Special Project (2024ZDZX0001-3), Sichuan Province Natural Science Foundation (Grant No.2024YFHZ0355), and the Science and Technology Development Fund, Macau SAR, under Grant 0193/2023/RIA3 and 0079/2025/AFJ. The authors would like to give special thanks to Dr. Wentao Feng for the workplace, computation power, and physical infrastructure support.
@@ D.3 Metrics
$\text{F1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
@@ E.1 Training Details for AIGC Benchmark Implementation
We resize images to 224×224 and apply only basic data augmentations, including flipping, brightness and contrast adjustment, compression, and Gaussian blur. All models are trained for 20 epochs using a cosine decay learning rate schedule, decreasing from 1e41e-4 to 1e51e-5. We use the ImageNet split of DiffusionForensics as the training set. For Synthbuster, we use a fully connected layer as the classifier.
@@ E.2 Training Details for Document Image Manipulation Localization Benchmark Models
Model FF++c40 $F$-DF $F$-F2F $F$-NT $F$-FS CDFv1 CDFv2 DFD DFDC DFDCP FaceShifter UADFV Average
@@ Appendix F Details of IFF-Protocol
Experiments on Recent Datasets.
We added experiments on the DF40 [74] and Chameleon [71] dataset, along with evaluations of two recent Deepfake SoTAs: Sia [60] (ECCV22) and Effort [73] (ICML25), and two recent AIGC SoTAs: FatFormer [33] (CVPR24) and CO-SPY [7] (CVPR25). Results are shown in Table 15.
Table 15: AUC performance on recent DF40 and Chameleon datasets.
Model DF40_CollabDiff DF40_deepfacelab DF40_heygen Chameleon
ConvNeXt 0.935 0.795 0.744 0.626
Capsule-Net 0.984 0.638 0.714 0.676
MVSS-Net 0.983 0.699 0.456 0.725
UnivFD 0.629 0.752 0.946 0.798
DTD 0.494 0.477 0.517 0.631
Sia 0.808 0.764 0.641 0.569
Effort 0.995 0.894 0.949 0.604
FatFormer 0.983 0.716 0.611 0.707
CO-SPY 0.890 0.792 0.656 0.767
Common features and conflicting patterns across domains.
We selected two IMDL models: MVSS-Net and IML-ViT, and used IFF-Protocol weights (where models are trained across-domain) as pretrained weights. These models were then trained on the IMDL task to investigate whether the artifacts learned across domains could benefit finetuning within a single domain. Results are shown in the Table 16.
Table 16: Common features and conflicting patterns across domains.
Method Coverage Columbia NIST16 CASIAv1 IMD2020 Avg.
MVSS-Net (Image-Net) 0.259 0.386 0.246 0.534 0.279 0.341
MVSS-Net (IFF-Protocol) 0.268 0.395 0.259 0.562 0.292 0.355
IML-ViT (Image-Net) 0.435 0.780 0.331 0.721 0.327 0.519
IML-ViT (IFF-Protocol) 0.427 0.767 0.279 0.715 0.351 0.508
@@ Appendix G Details of Experiments
Extractor Model FF++c40 $F$-DF $F$-F2F $F$-NT $F$-FS CDFv1 CDFv2 DFD DFDC DFDCP FaceShifter UADFV Mean
@@ G.2 Details for Extractor & Backbone in different tasks
We provide the performance differences of 6 backbones with and without 4 different feature extractors across the 4 domains. The table presents detailed results for each individual test dataset. They are Table 17 for IMDL, Table 18 for AIGC, Tabel 19 for Deepfake, and Table 20 for Document.
@@ Appendix H Computational Resources
[…]
@@ Appendix I Broader Impacts Discussion
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims reflect the paper’s contributions and scope, which we build the first unified benchmark for all-domain fake image detection and localization.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: We discuss the limitations in the Appendix.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: This paper does not include theory assumptions and proofs.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: We provide detailed experimental settings in this paper and open-source our code on GitHub.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: We provide detailed experimental settings in this paper and open-source our code on GitHub.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: We provide detailed experimental settings in this paper and open-source our code on GitHub.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [No]
  28. Justification: All experiments were conducted only once, but given the sufficiently large scale, the results can still serve as a valuable reference.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: We provide details of the experimental hardware setup in the Appendix.
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: All experiments in this paper are conducted using existing public datasets and comply with ethical standards.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: We discuss the proposed paper’s positive and negative societal impacts in the Appendix.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: This paper posed no such risks.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: All resources used in this paper are open-source and publicly available.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: We provide the detailed document along with the open-source code.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  57. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  63. Answer: [N/A]
  64. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.